OpenStudy (anonymous):

Solve. x^2 - 5x- 6 = 0

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

have you used the quadratic equation yet?

OpenStudy (anonymous):

yes and i got \[-1\pm \sqrt{49} /2\]

OpenStudy (anonymous):

maybe the quadratic equation was used wrong. it should be 5/2 +- sqrt(25-24)/2 = 5/2 +- 1/2

OpenStudy (anonymous):

should be 3 and 2